What is the coefficient of viscosity?
1
Expert's answer
2019-07-29T08:24:40-0400

Viscosity coefficient is defined in two ways:

1) dynamic viscosity of fluid show its internal resistance to flow due to shearing force;

2) kinematic viscosity is a ratio of dynamic viscosity to the density of fluid.


Dynamic viscosity is given by formula:


"η = \frac{F×h}{A×u}"

Dynamic viscosity is measured in


"\frac{N×s}{m^2}"



Kinematic viscosity is given by formula:


"k = \frac{ η }{ρ}"

where η is a dynamic viscosity, ρ is density of fluid

Kinematic viscosity is measured in



"\frac{m}{s^2}"